2x-3y equals -24
x+6y equals 18

Answer by ikleyn(52781)

.
2x-3y equals -24
x+6y equals 18
~~~~~~~~~~~~~~~~~~~~~~

2x - 3y = -24,   (1)
 x + 6y =  18.   (2)

From (2), express x = 18 - 6y and substitute it into (1). You will get a single equation for y

2*(18-6y) - 3y = -24.

Simplify and solve it.

From this point please complete your assignment yourself.